OpenStudy (anonymous):

The center of the following circle is located at (2, 4) x^2+y^2=4

OpenStudy (anonymous):

true or false ?

jimthompson5910 (jim_thompson5910):

Hint: x^2+y^2=4 is the same as (x-0)^2+(y-0)^2=2^2

OpenStudy (anonymous):

the center is in (0,0)